Problem 6-33 Accounting for depletion LO 6-8 Flannery Company engages in the exploration and development of many types of natural resources. In the last two years, the company has engaged in the following activities: Jan. 1, 2018 Purchased for $1,590,000 a silver mine estimated to contain 116,000 tons of silver ore. July 1, 2018 Purchased for $1,890,000 a tract of timber estimated to yield 1,720,000 board feet of lumber and the residual value of the land was estimated at $126,000. Feb. 1, 2019 Purchased for $2,950,000 a gold mine estimated to yield 68,000 tons of gold-veined ore. Aug. 1, 2019 Purchased oil reserves for $1,315,000. The reserves were estimated to contain 227,000 barrels of oil, of which 17,000 would be unprofitable to pump. Required (For all requirements, Round "per" values to 2 decimal places and final answers to the nearest whole dollar amount.) Determine the amount of depletion expense that would be recognized on the 2018 income statement for each of the two reserves, assuming 14,900 tons of silver were mined, and 570,000 board feet of lumber were cut. Determine the amount of depletion expense that would be recognized on the 2019 income statement for each of the four reserves, assuming 24,000 tons of silver are mined, 320,000 board feet of lumber are cut, 4,600 toms on gold ore is mined, and 52,000 barrels of oil are extracted. Prepare the portion of the December 31, 2019, balance sheet that reports natural resources.
Answer #1

Solution:

In this question, we are using the cost method of depletion.

Depletion formula    =          (Cost of Natural Resource – Salvage Value) / Estimated total capacity

(1) Computation of amount of depletion expense that would be recognised on 2018 income statement:

Silver Mine    = ($1,590,000 - $0)/116,000 = $13.71 depletion per tons of silver

Timber Land = ($1,890,000 - $126,000)/1,720,000 = $1.03 depletion per board feet

Depletion for the year 2018 will be:

For Silver Mine = $13.71 x 14,900 tons = $204,279

For Timber Lane = $1.03 x 570,000 board feet = $587,100

Total Depletion expense that would be recognised on 2018 income statement = $791,379

(2) Computation of amount of depletion expense that would be recognised on 2019 income statement:

Silver Mine    = ($1,590,000 - $0)/116,000 = $13.71 depletion per tons of silver

Timber Land = ($1,890,000 - $126,000)/1,720,000 = $1.03 depletion per board feet

Gold Mine = ($2,950,000 - $0)/68,000 = $43.38 depletion per tons of gold ore

Oil Reserve = ($1,315,000 - $0)/227,000 = $5.79 depletion per barrel

Depletion for the year 2018 will be:

For Silver Mine = $13.71 x 24,000 tons = $329,040

For Timber Lane = $1.03 x 320,000 board feet = $329,600

For Gold Mine = $43.38 x 4,600 tons = $199,548

For Oil Reserve = $5.79 x 52,000 barrels = $301,080

Total Depletion expense that would be recognised on 2019 income statement = $1,159,268

(3) Balance sheet as on 31st December 2019:

Balance Sheet as on 31st December 2019 Particulars Note No. Year End 2019 Year End 2018 EQUITY AND LIABILITIES ASSETS Non-cur

Notes to Accounts 1. Tangible Assets Silver Mine Opening Balance Less: Accumulated depletion Closing Balance (amount in $) Ye